# Pointwise convergence of ergodic averages with Möbius weight

### v1

@@ 5.3. A simple $L^{1}$ bound
We will also need a simple $L^{1}$ estimate for ergodic averages that follows from Hölder’s inequality.
@@
Let $k\in\mathbb{N}$. Let $(X,\nu)$ be a probability space, and let $T_{1},\ldots,T_{k}\colon X\to X$ be invertible measure-preserving maps. Let $g_{1},\ldots,g_{k}\colon\mathbb{Z}\to\mathbb{Z}$ be functions. Let $S\subset\mathbb{N}$ be a finite and nonempty. Also let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}\leq 1$. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$, we have
$\|\mathbb{E}_{n\in\mathcal{S}}f_{1}(T_{1}^{g_{1}(n)}x)\cdots f_{k}(T_{k}^{g_{k}(n)}x)\|_{L^{1}(X)}\leq\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)}.$
@@
Let $r\geq 1$ be such that $\frac{1}{r}=\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}$. Then, by Jensen’s and Hölder’s inequalities and the $T$-invariance of $\nu$, we have
$\|\mathbb{E}_{n\in\mathcal{S}}f_{1}(T_{1}^{g_{1}(n)}x)\cdots f_{k}(T_{k}^{g_{k}(n)}x)\|_{L^{1}(X)}$
$\quad\leq\|\mathbb{E}_{n\in\mathcal{S}}f_{1}(T_{1}^{g_{1}(n)}x)\cdots f_{k}(T_{k}^{g_{k}(n)}x)\|_{L^{r}(X)}$
$\quad\leq\left(\int_{X}\mathbb{E}_{n\in\mathcal{S}}|f_{1}(T_{1}^{g_{1}(n)}x)|^{q_{1}}\textnormal{ d}\nu(x)\right)^{1/q_{1}}\cdots\left(\int_{X}\mathbb{E}_{n\in\mathcal{S}}|f_{k}(T_{k}^{g_{k}(n)}x)|^{q_{k}}\textnormal{ d}\nu(x)\right)^{1/q_{k}}$
$\quad=\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)},$
@@ 6. Proofs of the pointwise ergodic theorems
where $\theta\colon\mathbb{N}\to\mathbb{C}$ is a function satisfying suitable hypotheses. Some of these hypotheses turn out to be trivial when $|\theta|\leq 1$ (so in particular for $\theta=\mu$), but the hypotheses are tailored so that the results are also applicable to a suitably normalised version of the von Mangoldt function (by the main results of [22][26]).
We start with Theorem 1.2. This theorem is an immediate consequence of Lemma 5.1(i),(iii) and the following theorem concerning the averages (6.1).
Theorem 6.1 (Pointwise convergence of some polynomial ergodic averages with nice weight).
Let $d,k,K\in\mathbb{N}$. Let $\theta\colon\mathbb{N}\to\mathbb{C}$ satisfy $|\theta(n)|\leq(\log n)^{K}$ for all $n\in\mathbb{N}$. Suppose also that, for any $M\geq 3$ and $B>0$, we have
(6.2) $\begin{cases}\|\theta\|_{U^{k+1}[M]}\ll_{B}(\log M)^{-B}\quad&\textnormal{if}\quad d=1\\ \|\theta\|_{u^{d+1}[M]}\ll_{B}(\log M)^{-B}\quad&\textnormal{if}\quad d\geq 2.\end{cases}$
Let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients satisfying $\deg P_{1}\leq\cdots\leq\deg P_{k}=d$. Suppose that one of the following holds:
@@
    $d\geq 2$ and $P_{1},\ldots,P_{k}$ have pairwise distinct degrees.
@@
Remark 6.2.
We briefly comment on what Theorem 6.1 tells us about the convergence of
(6.3) $\mathbb{E}_{n\in[N]}\theta(n)f_{1}(T^{n}x)f_{2}(T^{P(n)}x)$
for other bounded multiplicative functions $\theta$ than the Möbius function. Firstly, if we aim for convergence to 00, we should assume that $\theta$ is aperiodic (that is, it has mean 00 on every arithmetic progression). If we make no further assumption on $\theta$ than aperiodicity, even the $U^{1}[N]$ norm of $\theta$ could tend to zero arbitrarily slowly, so Theorem 6.1 cannot be applied in this case. But if we assume more strongly that $\theta$ satisfies the Siegel–Walfisz assumption (that is, $|\sum_{n\in\mathcal{P}}\theta(n)|\ll_{A}N/(\log N)^{A}$ for any arithmetic progression $\mathcal{P}\subset[N]$), and that $\theta$ has a combinatorial decomposition into type I and II sums, then it should be possible to verify the second bound in (6.2), leading to the convergence of (6.3) in the case $\deg P\geq 2$. But to obtain the first bound in (6.2) (which corresponds to $\deg P=1$ in (6.3)) we would need to impose an assumption stronger than the Siegel–Walfisz assumption in view of the best known $U^{3}$ inverse theorem [21, Theorem 7], which is quasipolynomial.
We shall reduce Theorem 6.1 to the following quantitative $L^{1}$ estimate.
@@
We are going to apply Lemma 5.4(i). Recall the notation (6.1). By Lemma 5.5 and the assumption $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$, for any $B>0$, $N\geq 3$ and $M\in[N,(1+(\log N)^{-B})N]$, we have
$(\log N)^{-K}\|A_{M}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})-A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})\|_{L^{1}(X)}$
$\quad\leq\int_{X}\left(\frac{1}{N}\sum_{M\leq n\leq N}\prod_{j=1}^{k}|f_{j}(T^{P_{j}(n)}x)|+\left(\frac{1}{M}-\frac{1}{N}\right)\sum_{n\leq N}\prod_{j=1}^{k}|f_{j}(T^{P_{j}(n)}x)|\right)\textnormal{ d}\nu(x)$
$\quad\ll(\log N)^{-B}\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)}.$
Hence, by Lemma 5.4(i), the conclusion of Theorem 6.1 follows from (6.4). ∎
@@
Reduction to the case of bounded functions. We claim that it suffices to prove Proposition 6.3 in the case $q_{1}=\cdots=q_{k}=\infty$. Suppose that this case has been proven. Let $1<q_{1},\ldots,q_{k}<\infty$ and $f_{i}\in L^{q_{i}}(X)$ for $i\in\{1,\ldots,k\}$ (we can assume that all the $q_{i}$ are <absent<$\infty$< ∞, since we have $\|f\|_{L^{q}(X)}\leq\|f\|_{L^{\infty}(X)}$ for any $q<\infty$ and any measurable $f$). Also let $A>0$ and let $C>0$ be large enough in terms of $A,K,q_{1},\ldots q_{k}$.
@@
Since $\|(\log N)^{-C}f_{j,N}^{(1)}\|_{L^{\infty}(X)}\leq 1$ for $j\in\{1,\ldots,k\}$, by the case $q_{1}=\cdots=q_{k}=\infty$ we have
@@
$\|A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})})\|_{L^{1}(X)}\ll_{C,q_{1},q_{2}}(\log N)^{K-cC}\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)}$
for some $c=c(q_{1},\ldots,q_{k})>0$, since $C$ can be taken to be large enough in terms of $A,K,q_{1},\ldots,q_{k}$ so that $K-cC\leq-A$.
Fix some $q_{i}^{\prime}\in(1,q_{i})$ such that $\frac{1}{q_{1}^{\prime}}+\cdots+\frac{1}{q_{k}^{\prime}}=1$, and let $c=\frac{q_{1}}{q_{1}^{\prime}}+\cdots+\frac{q_{k}}{q_{k}^{\prime}}-k$. Recalling the assumption $|\theta(n)|\leq(\log n)^{K}$ and using Lemma 5.5, we get
@@
$\quad\leq(\log N)^{K-cC}\|A_{N}^{P_{1},\ldots,P_{k}}(1;|f_{1,N}^{(i_{1})}|^{q_{1}/q_{1}^{\prime}},\ldots,|f_{k,N}^{(i_{k})}|^{q_{k}/q_{k}^{\prime}})\|_{L^{1}(X)}$
$\quad\ll(\log N)^{K-cC}\|f_{1}\|_{L^{q_{1}}(X)}^{q_{1}/q_{1}^{\prime}}\cdots\|f_{k}\|_{L^{q_{k}}(X)}^{q_{k}/q_{k}^{\prime}}.$
@@
$\int_{X}|A_{N}(x)|^{2}\textnormal{ d}\nu(x)\ll_{A,K,a,P}\|f_{1}\|_{L^{\infty}(X)}^{2}\cdots\|f_{k}\|_{L^{\infty}(X)}^{2}(\log N)^{-A}.$
The assumption $|\theta(n)|\leq(\log n)^{K}$ gives $\|A_{N}\|_{L^{\infty}(X)}\ll\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}(\log N)^{K}$, so it suffices to show that for any $\phi\in L^{\infty}(X)$ we have
(6.6) $\int_{X}\phi(x)A_{N}(x)\textnormal{ d}\nu(x)\ll_{A,K,a,P}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-A}.$

### v2

@@ 6.2. A simple $L^{1}$ bound
We also need a simple $L^{1}$ estimate for ergodic averages that follows from Hölder’s inequality.
@@
Let $k\in\mathbb{N}$. Let $(X,\nu)$ be a probability space, and let $T_{1},\ldots,T_{k}\colon X\to X$ be invertible measure-preserving maps. Let $g_{1},\ldots,g_{k}\colon\mathbb{Z}\to\mathbb{Z}$ be functions. Also let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}\leq 1$. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$, $\theta\colon\mathbb{N}\to\mathbb{C}$ and $N\geq 1$, we have
$\|\mathbb{E}_{n\leq N}\,\theta(n)\prod_{j=1}^{k}f_{j}(T_{j}^{g_{j}(n)}\cdot)\|_{L^{1}(X)}\leq\left(\mathbb{E}_{n\leq N}|\theta(n)|^{r}\right)^{1/r}\|f_{1}\|_{L^{q_{1}}(X)}\cdots\|f_{k}\|_{L^{q_{k}}(X)},$
where $1\leq r\leq\infty$ satisfies $11/r+1/q_{1}+\cdots+1/q_{k}=$ 1.
@@
By Hölder’s inequality and the $T$-invariance of $\nu$, we have
$\|\mathbb{E}_{n\leq N}\theta(n)\prod_{j=1}^{k}f_{j}(T_{j}^{g_{j}(n)}\cdot)\|_{L^{1}(X)}$
$\quad\leq\left(\mathbb{E}_{n\leq N}|\theta(n)|^{r}\right)^{1/r}\prod_{j=1}^{k}\left(\int_{X}\mathbb{E}_{n\in\mathcal{S}}|f_{j}(T_{j}^{g_{j}(n)}x)|^{q_{j}}\textnormal{ d}\nu(x)\right)^{1/q_{j}}$
$\quad=\left(\mathbb{E}_{n\leq N}|\theta(n)|^{r}\right)^{1/r}\prod_{j=1}^{k}\|f_{j}\|_{L^{q_{j}}(X)},$
@@ 7. Proofs of the pointwise ergodic theorems
where $\theta\colon\mathbb{N}\to\mathbb{C}$ is a function satisfying suitable uniformity norm estimates. We note the following result does not require $\theta$ to be multiplicative, and the hypotheses are satisfied also for example for $\theta$ being a suitably normalised version of the von Mangoldt function (by [25, Theorem 6]).
Theorem 7.1 (Pointwise convergence of polynomial ergodic averages with nice weight).
Let $d,k,K\in\mathbb{N}$. Let $\theta\colon\mathbb{N}\to\mathbb{C}$ satisfy
(7.2) $|\theta(n)|\leq(\log n)^{K}d(n)^{K}$
for all $n\geq 3$. Let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients satisfying $\deg P_{1}\leq\cdots\leq\deg P_{k}=d$. Suppose that one of the following holds:
@@
    We have
    $\|\theta\|_{u^{d+1}[M]}\ll_{B}(\log M)^{-B}$
    for any $M\geq 3$ and $B>0$, and the polynomials $P_{1},\ldots,P_{k}$ have pairwise distinct degrees.
@@
Let us first see how this theorem implies our main theorems.
Proof of Theorems 1.2 and 1.6 assuming Theorem 7.1.
Theorem 1.2 follows by taking $\theta=\mu$ and applying Lemma 5.1.
For proving Theorem 1.6, we first use Proposition 5.4 to obtain a decomposition $g=g_{1}+g_{2}$ with $\|g_{1}\|_{u^{k}[N]}\ll_{B,k}(\log N)^{-B}$, $\mathbb{E}_{n\leq N}|g_{2}(n)|^{r}\ll_{r}(\log N)^{-1+o(1)}$ and $|g_{1}|,|g_{2}|\leq|g|$. Applying Theorem 7.1, we reduce to showing that
$\lim_{N\to\infty}\frac{1}{N}\sum_{n\leq N}g_{2}(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
Since $11/q_{1}+\cdots+1/q_{k}<$ 1, we can find some $0<\varepsilon<\min_{1\leq j\leq k}(q_{j}-1)$ and $r\geq 1$ such that $11/r+1/(q_{1}-\varepsilon)+\cdots+1/(q_{k}-\varepsilon)=$ 1. Then by Hölder’s inequality we have
$\left|\mathbb{E}_{n\leq N}g_{2}(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)\right|$
$\quad\leq\left(\mathbb{E}_{n\leq N}|g_{2}(n)|^{r}\right)^{1/r}\prod_{j=1}^{k}\left(\mathbb{E}_{n\leq N}|f_{j}(T^{P_{j}(n)}x)|^{q_{j}-\varepsilon}\right)^{1/(q_{j}-\varepsilon)}.$
In view of the bound $\mathbb{E}_{n\leq N}|g_{2}(n)|^{r}\ll_{r}(\log N)^{-1+o(1)}$, it suffices to show for small enough $\delta>0$ that for all $1\leq j\leq k$ we have
(7.3) $\nu(\{x\in X\colon\limsup_{N\to\infty}\,(\log N)^{-\delta}\mathbb{E}_{n\leq N}|f_{j}(T^{P_{j}(n)}x)|^{q_{j}-\varepsilon}\geq 1\})=0.$
Using Markov’s inequality and Bourgain’s maximal inequality for polynomial ergodic averages (see [22, Theorem 1.8(iii)]) together with the fact that $f_{j}^{q_{j}-\varepsilon}\in L^{q_{j}/(q_{j}-\varepsilon)}(X)$, we see that for any $1\leq j\leq k$ we have
$\nu(\{x\in X\colon\sup_{N\in[2^{2^{m}},2^{2^{m+1}}]}\mathbb{E}_{n\leq N}|f_{j}(T^{P_{j}(n)}x)|^{q_{j}-\varepsilon}\geq 2^{(\delta/2)m}\})$
$\quad\leq(2^{-(\delta/2)m})^{-q_{j}/(q_{j}-\varepsilon)}\int_{X}\left(\sup_{N\geq 1}\mathbb{E}_{n\leq N}|f_{j}(T^{P_{j}(n)}x)|^{q_{j}-\varepsilon}\right)^{q_{j}/(q_{j}-\varepsilon)}\textnormal{ d}\nu(x)$
$\quad\ll 2^{-q_{j}\delta m/(2(q_{j}-\varepsilon))}\|f_{j}\|_{L^{q_{j}}(X)}^{q_{j}}.$
Since $\sum_{m\geq 1}2^{-cm}<\infty$ for any $c>0$, the claim (7.3) follows from the above estimate and the Borel–Cantelli lemma. ∎
We will reduce Theorem 7.1 to the following quantitative $L^{1}$ estimate.
@@
We are going to apply Lemma 6.1. Recall the notation (7.1). Let $r\geq 1$ satisfy $11/r+1/q_{1}+\cdots+1/q_{k}=$ 1. Applying first the triangle inequality, then Lemma 6.2 and finally (7.2) and Shiu’s bound, we see that for any $B>0$, $N\geq 3$ and $M\in[N,(1+(\log N)^{-(B/2-1)})N]$ we have
$\|A_{M}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})-A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1},\ldots,f_{k})\|_{L^{1}(X)}$
$\quad\leq\int_{X}\left(\frac{1}{N}\sum_{N\leq n\leq M}|\theta(n)|\prod_{j=1}^{k}|f_{j}(T^{P_{j}(n)}x)|+\left(\frac{1}{N}-\frac{1}{M}\right)\sum_{n\leq N}|\theta(n)|\prod_{j=1}^{k}|f_{j}(T^{P_{j}(n)}x)|\right)\textnormal{ d}\nu(x)$
$\quad\ll\left(\frac{M-N}{N}\left(\mathbb{E}_{N\leq n<M}|\theta(n)|^{r}\right)^{1/r}+\frac{(\log N)^{-(B/2-1)}}{N}\left(\mathbb{E}_{n\leq N}|\theta(n)|^{r}\right)^{1/r}\right)\prod_{j=1}^{k}\|f_{j}\|_{L^{q_{j}}(X)}$
$\ll(\log N)^{-B/2+1+K+(2^{Kr}-1)/r}\prod_{j=1}^{k}\|f_{j}\|_{L^{q_{j}}(X)}.$
If $B$ is large enough, the exponent of the logarithm above is at most $-0.49B$, say. Hence, by Lemma 6.1, the conclusion of Theorem 7.1 follows from (7.4). ∎
@@
Reduction to the case of bounded functions. We claim that it suffices to prove Proposition 7.2 in the case $q_{1}=\cdots=q_{k}=\infty$. Suppose that this case has been proven. Let $1<q_{1},\ldots,q_{k}<\infty$ and $f_{i}\in L^{q_{i}}(X)$ for $i\in\{1,\ldots,k\}$ (we can assume that all the $q_{i}$ are <absent<$\infty$< ∞, since we have $\|f\|_{L^{q}(X)}\leq\|f\|_{L^{\infty}(X)}$ for any $q<\infty$ and any measurable $f$). Also let $A>0$, and let $C>0$ be large enough in terms of $A,K,q_{1},\ldots q_{k}$.
@@
Since $\|(\log N)^{-C}f_{j,N}^{(1)}\|_{L^{\infty}(X)}\leq 1$ for $j\in\{1,\ldots,k\}$, by the case $q_{1}=\cdots=q_{k}=\infty$ of Proposition 7.2 we have
@@
$\|A_{N}^{P_{1},\ldots,P_{k}}(\theta;f_{1,N}^{(i_{1})},\ldots,f_{k,N}^{(i_{k})})\|_{L^{1}(X)}\ll_{C,q_{1},\ldots,q_{k}}(\log N)^{K+(2^{Kr}-1)/r-\delta C}\prod_{j=1}^{k}(1+\|f_{j}\|_{L^{q_{j}}(X)}^{1+\delta})$
for some $\delta=\delta(q_{1},\ldots,q_{k})>0$, since $C$ can be taken to be large enough in terms of $A,K,q_{1},\ldots,q_{k}$ so that $K(2^{Kr}-1)/r-\delta C\leq-A$.
For $(i_{1},\ldots,i_{k})\in[2]^{k}\setminus(1,\ldots,1)$, there is some $\ell$ such that $i_{\ell}=2$; for the sake of notation, assume that $\ell=1$. Fix some $\delta>0$ and $1\leq r<\infty$ such that $\frac{1}{r}+\frac{1+\delta}{q_{1}}+\frac{1}{q_{2}}+\cdots+\frac{1}{q_{k}}=1$. Using the triangle inequality, Lemma 6.2 and Shiu’s bound combined with the assumption $|\theta(n)|\leq(\log n)^{K}d(n)^{K}$, we obtain
@@
$\quad\leq(\log N)^{-\delta C}\|A_{N}^{P_{1},\ldots,P_{k}}(\theta;|f_{1,N}^{(i_{1})}|^{1+\delta},\ldots,|f_{k,N}^{(i_{k})}|)\|_{L^{1}(X)}$
$\quad\leq(\log N)^{-\delta C}\left(\mathbb{E}_{n\leq N}|\theta(n)|^{r}\right)^{1/r}\|f_{1}\|_{L^{q_{1}}(X)}^{1+\delta}\cdots\|f_{k}\|_{L^{q_{k}}(X)}$
$\quad\ll_{K,r}(\log N)^{K+(2^{Kr}-1)/r-\delta C}\|f_{1}\|_{L^{q_{1}}(X)}^{1+\delta}\cdots\|f_{k}\|_{L^{q_{k}}(X)}.$
@@
$\int_{X}|A_{N}(x)|^{2}\textnormal{ d}\nu(x)\ll_{A,K,P}\|f_{1}\|_{L^{\infty}(X)}^{2}\cdots\|f_{k}\|_{L^{\infty}(X)}^{2}(\log N)^{-A}.$
The assumption $|\theta(n)|\leq(\log n)^{K}d(n)^{K}$ gives
$\|A_{N}\|_{L^{\infty}(X)}\ll(\log N)^{K+2^{K}-1}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)},$
so it suffices to show that for any $\phi\in L^{\infty}(X)$ we have
(7.6) $\int_{X}\phi(x)A_{N}(x)\textnormal{ d}\nu(x)\ll_{A,K,P}\|f_{1}\|_{L^{\infty}(X)}\cdots\|f_{k}\|_{L^{\infty}(X)}\|\phi\|_{L^{\infty}(X)}(\log N)^{-A}.$